# Yang-Lee Quantum Criticality in Various Dimensions

### v3

@@ 1 Introduction and Summary
In $d=2$, the exact Ising critical exponents have been known for many years thanks to the Onsager solution [2]. The critical 2D Ising model corresponds to the simplest unitary Belavin-Polyakov-Zamolodchikov (BPZ) conformal minimal model $M(3,4)$ [3]. More generally, the critical $d$-dimensional Ising model is described by the continuum theory of a massless scalar field $\phi$ with the interaction term $\sim\phi^{4}$. The problem of finding the critical exponents is then mapped to calculation of the scaling dimensions of operators $\phi$, $\phi^{2}$, $\phi^{4}$, etc. In dimensions slightly below $4$, the IR fixed point is weakly coupled, and the Wilson-Fisher $4-\epsilon$ expansion [4] provides a useful method for approximating the $d$ dependence of operator dimensions. This has led to estimates of the 3D Ising critical exponents that are in very good agreement with experiments and Monte Carlo simulations. A significant improvement in accuracy was provided by the numerical conformal bootstrap [5, 6], which is based on the conformal symmetry of critical fluctuations [7] (for reviews, see [8, 9, 10, 11]). Furthermore, the numerical bootstrap method has been applied to the operator scaling dimensions in the entire range $42\leq d<$ 4 [12, 13], producing excellent agreement with resummed $4-\epsilon$ expansions for the $\phi^{4}$ theory.
@@
A fascinating feature of the $d$-dimensional Ising model is the appearance of zeros of the partition function $Z$ at imaginary values of the magnetic field [35, 36]. In the paramagnetic phase $T>T_{c}$, these Yang-Lee (YL) zeros become dense near the edge value $h_{\rm crit}(T)$. As shown by Kortman and Griffiths [37], the density of zeros has the scaling form $\rho(h)\sim(h-h_{\rm crit}(T))^{\sigma}$, and $\sigma$ is negative in $d=2$ but positive in $d=3$. In an important further development, Michael Fisher [38] realized that this critical behavior is described by the Euclidean field theory with the imaginary potential $\sim i\phi^{3}$. In this continuum approach, the critical exponent is $\sigma=\frac{\Delta_{\phi}}{d-\Delta_{\phi}}$, where $\Delta_{\phi}$ is the scaling dimension of the real scalar field $\phi$. The field theory description of the YL universality class implies that its upper critical dimension is $6$. The lower critical dimension is $1$, and the exact solution there gives $\sigma=-\frac{1}{2}$ [35, 36]. Therefore, an interesting problem is to determine the $d$ dependence of the critical exponents in the range $61<d<$ 6. One approach, analogous to the Wilson-Fisher $4-\epsilon$ expansion for the Ising universality class, is to develop the $6-\epsilon$ expansion for the massless $i\phi^{3}$ field theory. This approach, in conjunction with the exact result in $d=1$, was used by Fisher [38] to estimate the $d$-dependence of $\sigma$ for the YL universality class. Soon after the advent of the BPZ minimal models [3], Cardy [39] realized that the critical YL model is also exactly solvable in $d=2$, where it corresponds to the simplest non-unitary minimal model $M(2,5)$ (for a recent review, see [40]). The field $\phi$ corresponds to the Virasoro primary $\Phi_{1,2}$ of holomorphic dimension $h_{1,2}=-\frac{1}{5}$. This means that $\Delta_{\phi}=-\frac{2}{5}$ and $\sigma=-\frac{1}{6}$ in $d=2$.
@@
In the range $62<d<$ 6, no similar exact results are known. In this paper, we will use a combination of $6-\epsilon$ expansions and numerical diagonalizations to approximately determine the scaling dimensions of various operators. The $6-\epsilon$ expansions are available to $O(\epsilon^{5})$ [41, 42]. [^1] After the original version of this paper appeared, the $6$-loop beta function and anomalous dimension became available [43]; they determine the $O(\epsilon^{6})$ terms in $\Delta_{\phi}$ and $\Delta_{\phi^{3}}$. Their resummations have been extrapolated to $d=2$, yielding good agreement with $M(2,5)$. However, for the higher dimension operators this method can exhibit significant uncertainties. The numerical boostrap approach is complicated by the fact that the YL model is non-unitary, so that the standard methods based on positivity cannot be used. A different bootstrap approach to the YL model was used in [44, 45, 46], but it does not appear to be as precise as the numerical bootstrap based on positivity. We will instead follow the successes of the fuzzy sphere [14] and the icosahedron[17] [^2] The icosahedron has also been used to discretize the 2D classical Ising model on a sphere [47, 48] approaches to the 3D Ising model, adapting them to the YL case where a non-Hermitian Hamiltonian has to be tuned to quantum criticality. We will proceed analogously to the quantum spin chain approach which has proven to be successful for describing the 2D YL model. Here an imaginary longitudinal magnetic field $h_{z}$ was introduced and tuned to a critical value from below [49, 50, 51]. Even though the spin chain Hamiltonian is complex, the energy levels are real for $h_{z}<h_{z}^{\rm crit}$ due to the $\mathcal{PT}$ symmetry [51].
@@ 2 Minimal model $M(2,5)$ and its field theory description
,    I $\phi$ $\phi$ I
Table 1: $M(2,5)$ model with the central charge $c=-22/5$.
@@
denote the spin $4$ quasi-primary state [58]. A related spin 00 quasiprimary is $\Xi=Q\bar{Q}|\phi\rangle$ [59].
@@ 3 Spin chain for 2D Yang-Lee model
We set $J=1$ and assume that the system is in paramagnetic (disorder) phase, so the transverse magnetic field $h_{x}>h_{x}^{\textrm{crit}}$, where $h_{x}^{\textrm{crit}}/J=1$ is the critical value for the 2D Ising quantum model. It corresponds to the $M(3,4)$ model, whose Kac table is well known to be
@@
The Yang-Lee critical point is reached by tuning the imaginary magnetic field $ih_{z}$ to the critical point $ih_{z}^{\textrm{crit}}$. The Hamiltonian (3.1) was studied numerically long ago [52, 49, 50] and also more recently in [71, 72]. The qualitative phase structure is shown in Fig. 4; we will see that it applies also in higher dimensions.
@@
Figure 5: Energy gaps of the Hamiltonian (3.1) for $N=24$ along the path $(h_{x},ih_{z})=(0,0)\to(2,0)\to(2,i0.4)$, with $J=1$. We identified the lower energy levels with their corresponding operators in the 2D Ising and Yang-Lee CFTs. The first dashed lines is the critical point $h_{x}^{\textrm{crit}}=1$ for the 2D Ising model and the second dashed line is the pseudo-critical point $ih_{z}^{*}$ for the 2D Yang-Lee model, obtained fixing $r_{T}=2$.
@@
There are various examples of these criteria: Finite-Size-Scaling (FSS) [73, 74, 75], extrapolation of merging points [76, 77], CFT ratios of energy gaps [53], etc. We define the pseudo-critical point as the one where the following the criterion is satisfied:
$r_{T}=\frac{E_{T}-E_{\textrm{I}}}{E_{\partial\phi}-E_{\phi}}=2,$ (3.4)
@@
Figure 6: 3D Platonic solids $(N,E)$: Cube$(8,12)$, Icosahedron$(12,30)$, Dodecahedron$(20,30)$.
[…]
$h_{x}$ $5$ $10$ $15$ $20$ $25$ $30$ $35$ $40$ $45$ $50$
[…]
@@
$h_{x}$ $5$ $10$ $15$ $20$ $25$ $30$ $35$ $40$ $45$ $50$
[…]
@@ 5.1 Fuzzy sphere approach to Yang-Lee quantum criticality
×$\times\bigg{[}(\mathbf{c}_{m_{1}}^{\dagger}\mathbf{c}_{m_{4}})(\mathbf{c}_{m_{2}}^{\dagger}\mathbf{c}_{m_{3}})-(\mathbf{c}_{m_{1}}^{\dagger}\sigma^{z}\mathbf{c}_{m_{4}})(\mathbf{c}_{m_{2}}^{\dagger}\sigma^{z}\mathbf{c}_{m_{3}})\bigg{]}$
@@
is parametrized by Haldane pseudo-potential $V_{\ell}$. In our work we truncate the potential to $V_{0}$ and $V_{1}$, which is equivalent to the derivative expansion truncated to order $\nabla^{2}$. The overall normalization is fixed with $V_{1}=1$ and $V_{0}$ is a tunable parameter of the model. The LLL-projected Hamiltonian has $(4s+2)$ levels. We fill a the total number of $N=2s+1$ fermions in these levels. Under half-filling the charge degrees of freedom are gapped, and also the particle-hole symmetry is preserved. The total number of spins $N$ can be treated as a UV cutoff, which can be most clearly seen from the fact that the commutator $\{\psi(\Omega),\psi(\Omega^{\prime})\}$ approaches a delta function of size $\sqrt{1/N}$. This allows us to identify the radius of the sphere as $R\sim\sqrt{N}$.
@@
[…]
Figure 8: Operator flow of first spin 0 (red), spin 2 (blue) and spin 4 (purple) operators from critical Ising model ($h_{x}^{*}=3.16$) to operators in YL model with the same spin, following the trajectory depicted in Fig. 4. The YL pseudo-critical point is located at $ih^{*}_{z}\sim 4.5$ for the parameters $N=12$, $V_{0}=4.75$, and $h_{x}=10$. We see that operators that come from different $\mathbb{Z}_{2}$-Ising charge sectors, but the same spin, merge close to $ih_{z}^{*}$.
@@
Finally we notice that operators from different sectors of the Ising-$\mathbb{Z}_{2}$ symmetry and equal spin on the left ($ih_{z}=0$) merge in the Yang-Lee side on the right ($ih_{z}\neq 0$), where such symmetry is explicitly broken, e.g. I merges with $\sigma$ and $\Box\varepsilon$ merges with $\Box\sigma$ at some $ih_{z}$ value.

### v4

@@ 1 Introduction and Summary
In $d=2$, the exact Ising critical exponents have been known for many years thanks to the Onsager solution [2]. The critical 2D Ising model corresponds to the simplest unitary Belavin-Polyakov-Zamolodchikov (BPZ) conformal minimal model $M(3,4)$ [3]. More generally, the critical $d$-dimensional Ising model is described by the continuum theory of a massless scalar field $\phi$ with the interaction term $\sim\phi^{4}$. The problem of finding the critical exponents is then mapped to calculation of the scaling dimensions of operators $\phi$, $\phi^{2}$, $\phi^{4}$, etc. In dimensions slightly below $4$, the IR fixed point is weakly coupled, and the Wilson-Fisher $4-\epsilon$ expansion [4] provides a useful method for approximating the $d$ dependence of operator dimensions. This has led to estimates of the 3D Ising critical exponents that are in very good agreement with experiments and Monte Carlo simulations. A significant improvement in accuracy was provided by the numerical conformal bootstrap [5, 6], which is based on the conformal symmetry of critical fluctuations [7] (for reviews, see [8, 9, 10, 11]). Furthermore, the numerical bootstrap method has been applied to the operator scaling dimensions in the entire range $2\leq d<4$ [12, 13], producing excellent agreement with resummed $4-\epsilon$ expansions for the $\phi^{4}$ theory.
@@
A fascinating feature of the $d$-dimensional Ising model is the appearance of zeros of the partition function $Z$ at imaginary values of the magnetic field [35, 36]. In the paramagnetic phase $T>T_{c}$, these Yang-Lee (YL) zeros become dense near the edge value $h_{\rm crit}(T)$. As shown by Kortman and Griffiths [37], the density of zeros has the scaling form $\rho(h)\sim(h-h_{\rm crit}(T))^{\sigma}$, and $\sigma$ is negative in $d=2$ but positive in $d=3$. In an important further development, Michael Fisher [38] realized that this critical behavior is described by the Euclidean field theory with the imaginary potential $\sim i\phi^{3}$. In this continuum approach, the critical exponent is $\sigma=\frac{\Delta_{\phi}}{d-\Delta_{\phi}}$, where $\Delta_{\phi}$ is the scaling dimension of the real scalar field $\phi$. The field theory description of the YL universality class implies that its upper critical dimension is $6$. The lower critical dimension is $1$, and the exact solution there gives $\sigma=-\frac{1}{2}$ [35, 36]. Therefore, an interesting problem is to determine the $d$ dependence of the critical exponents in the range $1<d<6$. One approach, analogous to the Wilson-Fisher $4-\epsilon$ expansion for the Ising universality class, is to develop the $6-\epsilon$ expansion for the massless $i\phi^{3}$ field theory. This approach, in conjunction with the exact result in $d=1$, was used by Fisher [38] to estimate the $d$-dependence of $\sigma$ for the YL universality class. Soon after the advent of the BPZ minimal models [3], Cardy [39] realized that the critical YL model is also exactly solvable in $d=2$, where it corresponds to the simplest non-unitary minimal model $M(2,5)$ (for a recent review, see [40]). The field $\phi$ corresponds to the Virasoro primary $\Phi_{1,2}$ of holomorphic dimension $h_{1,2}=-\frac{1}{5}$. This means that $\Delta_{\phi}=-\frac{2}{5}$ and $\sigma=-\frac{1}{6}$ in $d=2$.
@@
In the range $2<d<6$, no similar exact results are known. In this paper, we will use a combination of $6-\epsilon$ expansions and numerical diagonalizations to approximately determine the scaling dimensions of various operators. The $6-\epsilon$ expansions are available to $O(\epsilon^{5})$ [41, 42]. [^1] After the original version of this paper appeared, the $6$-loop beta function and anomalous dimension became available [43]; they determine the $O(\epsilon^{6})$ terms in $\Delta_{\phi}$ and $\Delta_{\phi^{3}}$. Their resummations have been extrapolated to $d=2$, yielding good agreement with $M(2,5)$. However, for the higher dimension operators this method can exhibit significant uncertainties. The numerical boostrap approach is complicated by the fact that the YL model is non-unitary, so that the standard methods based on positivity cannot be used. A different bootstrap approach to the YL model was used in [44, 45, 46], but it does not appear to be as precise as the numerical bootstrap based on positivity. We will instead follow the successes of the fuzzy sphere [14] and the icosahedron[17] [^2] The icosahedron has also been used to discretize the 2D classical Ising model on a sphere [47, 48] approaches to the 3D Ising model, adapting them to the YL case where a non-Hermitian Hamiltonian has to be tuned to quantum criticality. We will proceed analogously to the quantum spin chain approach which has proven to be successful for describing the 2D YL model. Here an imaginary longitudinal magnetic field $h_{z}$ was introduced and tuned to a critical value from below [49, 50, 51]. [^3] More recent studies of non-Hermitian spin chain Hamiltonians were carried out in the context of $Q$-state Potts models with $Q=5$ [52, 53]. Let us also note that the YL universality class is related to the $Q\rightarrow\infty$ limit of the Potts model [54]. Even though the spin chain Hamiltonian is complex, the energy levels are real for $h_{z}<h_{z}^{\rm crit}$ due to the $\mathcal{PT}$ symmetry [51].
@@ 2 Minimal model $M(2,5)$ and its field theory description
,    I $\phi$ $\phi$ I ,
Table 1: $M(2,5)$ model with the central charge c=22/5c=-22/5.
@@
denote the spin $4$ quasi-primary state [61]. A related spin 0 quasiprimary is $\Xi=Q\bar{Q}|\phi\rangle$ [62].
@@ 3 Spin chain for 2D Yang-Lee model
We set $J=1$ and assume that the system is in paramagnetic (disordered) phase, so the transverse magnetic field $h_{x}>h_{x}^{\textrm{crit}}$, where $h_{x}^{\textrm{crit}}/J=1$ is the critical value for the 2D Ising quantum model. It corresponds to the $M(3,4)$ model, whose Kac table is well known to be
@@
The Yang-Lee critical point is reached by tuning the imaginary magnetic field $ih_{z}$ to the critical point $ih_{z}^{\textrm{crit}}$. At this critical point, some pairs of energy levels merge, and beyond it they become complex conjugates of each other. Since the Hamiltonian (3.1) is symmetric, $H_{\textrm{YL}}^{T}=H_{\textrm{YL}}$, it is pseudo-Hermitian: $H_{\textrm{YL}}=\mathcal{P}H_{\textrm{YL}}^{{\dagger}}\mathcal{P}$. In this case, each eigenstate can be assigned an appropriately defined eigenvalue of $\mathcal{PT}$, which is either $1$ or 1-1; it is also known as the Krein signature [74, 75]. For $h_{z}=0$, the eigenvalue of $\mathcal{PT}$ coincides with the $\mathbbm{Z}_{2}$ eigenvalue of the $\mathcal{P}$ operator. For two states to merge, they must carry opposite eigenvalues of $\mathcal{PT}$ [74, 75, 76].
The Hamiltonian (3.1) was studied numerically long ago [55, 49, 50] and also more recently in [77, 78]. Our numerical results, shown in Figure 4, are in agreement with these calculations and extend them. For a finite $N$, mergers of different pairs of levels, which are called “exceptional points” [79], occur at somewhat different values of $h_{z}$. However, they are expected to approach the same critical value as $N\rightarrow\infty$. The qualitative phase structure in the thermodynamic limit is shown in Fig. 5; we will see that it applies also in higher dimensions.
Figure 4 also shows that, as $h_{z}$ is increased beyond the merger point, there may be another transition, after which the energy levels become real again. Two energy levels becoming complex conjugates in a small region of parameter space is a common phenomenon in $\mathcal{PT}$-symmetric systems; it occurs due to a small imaginary off-diagonal element in a non-Hermitian Hamiltonian [79]. If the off-diagonal element is very small, the evolution of the two energy levels resembles their crossing. For the pseudo-Hermitian Hamiltonians, such as (3.1), a pair of energy levels can “cross” only if they have opposite eigenvalues of $\mathcal{PT}$ (Krein signatures) [74, 75, 76]. For a Hermitian Hamiltonian, a small off-diagonal element instead produces an avoided crossing of two energy levels.
Figure 4: Energy levels of the Hamiltonian (3.1) for $N=24$ along the path $(h_{x},ih_{z})=(0,0)\to(2,0)\to(2,i0.4)$, with $J=1$. Beyond a merger point, two energy levels become complex conjugate, and we plot their real part (sometimes another transition is visible where the two energies become real again). We identify the lower energy levels with their corresponding operators in the 2D Ising and Yang-Lee CFTs. The first dashed lines is the critical point $h_{x}^{\textrm{crit}}=1$ for the 2D Ising model and the second dashed line is the pseudo-critical point $ih_{z}^{*}$ for the 2D Yang-Lee model, obtained fixing $r_{T}=2$.
@@
There are various examples of these criteria: Finite-Size-Scaling (FSS) [80, 81, 82], extrapolation of merging points [83, 84], CFT ratios of energy gaps [56], etc. We define the pseudo-critical point as the point at which the following criterion is satisfied:
$r_{T}=\frac{E_{T}-E_{\textrm{I}}}{E_{\partial\phi}-E_{\phi}}=2\,.$ (3.4)
@@
Figure 6: 3D Platonic solids $(N,E)$: Cube(8,12)(8,12), Icosahedron(12,30)(12,30), Dodecahedron(20,30)(20,30).
These three Platonic solids have different ratio of number of vertices $N$ and edges $E$, therefore if we would like to compare their properties between each other we have to adjust couplings $J$ or $h_{x}$ accordingly. We keep $h_{x}$ the same for all these Platonic solids and adjust only the coupling $J$. We set $J_{\textrm{dodecahedron}}=1$ and adjust other solids’ $J$ couplings assuming that $JE/N=\textrm{const.}$. Therefore we find $J_{\textrm{cube}}=1$ and $J_{\textrm{icosahedron}}=3/5$. In the Table 6 we list values for the merger point $ih_{z}^{\textrm{merg}}$ of two lowest states: $|\textrm{I}\rangle$ and $|\phi\rangle$, which correspond to the lowest energy levels $E_{0}^{(\textbf{1})}$ and $E_{1}^{(\textbf{1})}$ in the singlet representation 1 of 3D symmetry point groups. Below we use notation $E_{n}^{(\textbf{r})}$ for the energy level $n$ in irrep r of the point group.
$h_{x}$ 55 $10$ $15$ 2020 $25$ $30$ $35$ $40$ $45$ $50$
$\textrm{C:}\;h_{z}^{\textrm{merg}}$ 1.513 5.173 9.283 13.591 18.016 22.517 27.074 31.674 36.307 40.968
$\textrm{I:}\;\;\,h_{z}^{\textrm{merg}}$ 1.339 4.926 8.991 13.265 17.661 22.138 26.673 31.253 35.868 40.513
$\textrm{D:}\;h_{z}^{\textrm{merg}}$ 1.387 5.020 9.112 13.406 17.817 22.308 26.855 31.438 36.071 40.723
@@
$h_{x}$ 55 $10$ $15$ 2020 $25$ $30$ $35$ $40$ $45$ $50$
$\textrm{C:}\;\Delta_{\phi}$ 0.424 0.346 0.323 0.312 0.305 0.300 0.297 0.294 0.292 0.290
$\textrm{I:}\;\;\Delta_{\phi}$ 0.385 0.306 0.283 0.272 0.265 0.260 0.257 0.254 0.252 0.250
$\textrm{D:}\;\Delta_{\phi}$ 0.289 0.244 0.233 0.228 0.224 0.222 0.221 0.220 0.219 0.218
@@
The results of this section show that the Platonic solid regulators of the sphere give a good indication of the YL quantum criticality, in spite of the fact that there are only 3 possible numbers of lattice sites. However, the determination of scaling dimensions is not as precise as in the fuzzy sphere method that we will use in the next section. In order to improve the Platonic solid approach to YL criticality, one can use the conformal perturbation theory, which was shown to work well for the Platonic solid approach to the 3D Ising model [17, 18]. We leave this for future work.
@@ 5.1 Fuzzy sphere approach to Yang-Lee quantum criticality
$\times\bigg[(\mathbf{c}_{m_{1}}^{\dagger}\mathbf{c}_{m_{4}})(\mathbf{c}_{m_{2}}^{\dagger}\mathbf{c}_{m_{3}})-(\mathbf{c}_{m_{1}}^{\dagger}\sigma^{z}\mathbf{c}_{m_{4}})(\mathbf{c}_{m_{2}}^{\dagger}\sigma^{z}\mathbf{c}_{m_{3}})\bigg]$
@@
is parametrized by the Haldane pseudo-potentials $V_{l}$. In our work we truncate the potential to $V_{0}$ and $V_{1}$, which is equivalent to the derivative expansion truncated to order $\nabla^{2}$. The overall normalization is fixed with $V_{1}=1$ and $V_{0}$ is a tunable parameter of the model. The LLL-projected Hamiltonian has $(4s+2)$ levels. We fill the total number of $N=2s+1$ fermions in these levels. Under half-filling the charge degrees of freedom are gapped, and also the particle-hole symmetry is preserved. The total number of spins $N$ can be treated as a UV cutoff, which can be most clearly seen from the fact that the commutator $\{\psi(\Omega),\psi(\Omega^{\prime})\}$ approaches a delta function of size $\sqrt{1/N}$. This allows us to identify the radius of the sphere as $R\sim\sqrt{N}$.
@@
We expect the phase diagram of the Hamiltonian (5.5) to also be Fig. 5. Similar to the 2D case, there is a phase transition when we take the Ising model in the paramagnetic phase and turn on an imaginary longitudinal magnetic field $h_{z}$. To show the qualitative behavior of the phase transition, it is helpful to identify the correspondence of low-lying eigenstates between Ising criticality and YL criticality. We scan coupling constants $h_{x}$ and $ih_{z}$ along the red-dashed trajectory in Fig. 5 and keep track of a number of eigenvalues. The Ising criticality is reported in [14] to occur at couplings $V_{0}=4.75$ and $h_{x}=3.16$. We hold $V_{0}$ fixed throughout the trajectory. We gradually increase $h_{x}$ from 0 to $10$, then increase $h_{z}$ from 0 to $h_{z}\approx 5.2$. The spectrum along this trajectory is shown in Fig. 8, we can identify the tower of operators that belong to the 3D YL universality class.
Figure 8: Energy gaps of the fuzzy sphere Hamiltonian (5.5) along the path $(h_{x},ih_{z})=(0,0)\to(10,0)\to(10,i5.2)$, where $N=12$, $V_{0}=4.75$. We identified the lower energy levels with their corresponding operators in the Ising and Yang-Lee CFTs. The dashed lines are the Ising critical point $h_{x}^{\textrm{crit}}\approx 3.16$ and the Yang-Lee pseudo-critical point $ih^{*}_{z}\approx 4.5$ obtained fixing $r_{T}=3$. We notice that operators with the same spin that merge near $ih_{z}^{*}$ always come from different $\mathbb{Z}_{2}$ Ising sectors, hence they have opposite eigenvalues of $\mathcal{PT}$.
@@
Finally we notice that operators from different sectors of the Ising-$\mathbb{Z}_{2}$ symmetry and equal spin on the left ($ih_{z}=0$) merge on the Yang-Lee side on the right ($ih_{z}\neq 0$), where such symmetry is explicitly broken, e.g. I merges with $\sigma$ and $\Box\varepsilon$ merges with $\Box\sigma$ at some $ih_{z}$ value. This is consistent with the general result, reviewed in section 3, that a pair of levels can merge only if they have opposite eigenvalues of $\mathcal{PT}$.